OpenStudy (anonymous):

the line with a slope of -1/4 and a y intercept of -5/4

OpenStudy (anonymous):

General Form of Straight Line: y=mx+c where m is the slope and c is the y-intercept. \(y=\frac{-1}{4}x-\frac{5}{4}\)

OpenStudy (anonymous):

y = -x/4 -5/4